# Ashish Saxena 11    Tanvir Kaur 11    Kaushik Mondal 11

### v2

@@ 2 Impossibility Results and Lower Bound
(Impossibility without 1-hop visibility) It is impossible to solve $k$-Dispersion on TVG for any $n\geq 6$ considering global communication but without $1$-hop visibility even if agents are initially co-located. This impossibility result also holds if agents have infinite memory and knowledge of $n$.
Proof
We prove this impossibility on TVG ring $\mathcal{R}$. Let $k=pn$ many agents be co-located at some node $v$ of ring $\mathcal{R}$, where $p$ is any positive integer. If the agents have zero-hop visibility, then the choice of movement of each agent depends only on the information present at the current node and the information attained via global communication. Since the algorithm is deterministic, the adversary can remove an edge from $G$ based on the movement of the agents. Let the agents follow algorithm $\mathcal{A}$ to achieve dispersion. In a given round, if the movement of agents does not lead to dispersion, the adversary does not need to do anything. If the movement of agents leads to dispersion, the adversary has to intervene. Assume that at the end of round $r$, the agents successfully achieve dispersion. It means that the dispersion is not achieved at the beginning of round $r$, and after the CCM cycle at round $r$, the agents are in the dispersed configuration. Since $k=pn$, each node of $\mathcal{R}$ should have $p$ many agents in the dispersed configuration. At the beginning of round $r$, there is a node $u$ which has at least $p+1$ many agents, and there is a node $v$ which has at most $p-1$ agents. Let $u_{1}$ and $u_{2}$ be two neighbours of node $u$ in ring $\mathcal{R}$. Based on the agent’s movement from node $u$, we give the strategy for the adversary at the beginning of round $r$ such that the dispersed configuration is not achieved at the end of round $r$. A high-level idea behind the impossibility proof is given in Figure 2. We divide the cases as follows.
  • Case 1. According to $\mathcal{A}$, if all the agents at node $u$ stay after $CCM$, then the adversary does not delete any edge at the beginning of round $r$. Since the agents of node $u$ do not move, therefore, at node $u$, there are at least $p+1$ agents at the end of round $r$. Hence, the dispersed configuration is not achieved at the end of round $r$.
  • Case 2. According to $\mathcal{A}$, at least one agent at node $u$ move to $u_{1}$ or $u_{2}$ after $CCM$ at round $r$. W.l.o.g., at least one agent at node $u$ moves to node $u_{1}$, and no agents move from $u$ to $u_{2}$. Let $x$ agents move to node $u_{1}$. The adversary deletes edge $(u,\,u_{1})$ at the beginning of round $r$. Since agents have 0-hop visibility, they can not understand the deletion of the edge at round $r$. Since edge $(u,u_{1})$ has been deleted, the movement of agents will be unsuccessful. In this case, at the end of round $r$, at node $u$, there will be at least $p+1$ agents. Hence, the dispersed configuration is not achieved at the end of round $r$.
  • Case 3. Let $z$ agents be at node $u$ at the beginning of round $r$, where $z>p$. According to $\mathcal{A}$ at round $r$, at least one agent at node $u$ moves to $u_{1}$, and at least one agent at node $u$ moves to node $u_{2}$. Let $x_{1}$ agents from node $u$ move to node $u_{1}$, and $y_{1}$ agents from node $u$ move to node $u_{2}$. Note that $x_{1},\,y_{1}\geq 1$. Let $x_{2}$ agents from node $u_{1}$ move to node $u$, and $y_{2}$ agents from node $u_{2}$ move to node $u$, where $x_{2},\,y_{2}\geq 0$. If the adversary does not delete any edge, then at the end of round $r$, there are $(z-x_{1}-y_{1}+x_{2}+y_{2})=p$ agents at node $u$. The adversary knows the precomputed values of $x_{1},x_{2},y_{1},y_{2}$ and accordingly deletes an edge. In this case, we divide this case into two sub-cases.
    • Sub-case $\bm{x_{1}=x_{2}}$ and $\bm{y_{1}=y_{2}:}$ At the end of round $r$, there are $z$ agents at node $u$, and $z>p$. Therefore, the dispersion configuration is not achieved at the end of round $r$.
    • Sub-case $\bm{x_{1}\neq x_{2}}$ or $\bm{y_{1}\neq y_{2}:}$ W.l.o.g., let $x_{1}\neq x_{2}$. In this case, the adversary deletes the edge $(u,\,u_{1})$ at the beginning of round $r$. Therefore, the movement of agents from node $u$ to $u_{1}$ and node $u_{1}$ to $u$ becomes unsuccessful. Since agents are unaware of the deletion of edge, the movement of agents from node $u$ to $u_{2}$ and node $u_{2}$ to $u$ becomes successful. Therefore, the number of agents at node $u$ is $z-y_{1}+y_{2}$. To achieve the dispersed configuration at the end of round $r$, the value of $z-y_{1}+y_{2}$ must be $p$. Therefore, $z-y_{1}+y_{2}=z-x_{1}-y_{1}+x_{2}+y_{2}\implies x_{1}=x_{2}$, and it is a contradiction due to $x_{1}\neq x_{2}$. Therefore, the value of $z-y_{1}+y_{2}$ is either less than $p$ or greater than $p$. Hence, the dispersion configuration is not achieved at the end of round $r$.
This completes the proof. ∎
Remark 1
It is important to note that the proof of both impossibility results is valid when $k=pn$, where $p$ is some constant, i.e., agents need 1-hop visibility and global communication to solve $pn$-Dispersion on TVG. In Section 3, we provide an algorithm to solve $(n+1)$-Dispersion on TVG with local communication and zero-hop visibility, and extend this idea in Section 4.1 to solve $(pn+1)$-Dispersion on TVG. The aforementioned proofs do not hold when $k\in[pn+2,\,(p+1)(n-1)]$, where $p\in\mathbb{N}$. Therefore, it is an open question whether 1-hop visibility and global communication are necessary to solve $k$-Dispersion on TVG for $k\in[pn+2,\,(p+1)(n-1)]$, where $p\in\mathbb{N}$. Also, it remains open to come up with the necessary conditions to solve $k$-Dispersion on TVG for $k<n$ given the algorithm of [16] that works with global communication and 1-hop visibility is a valid solution.
Remark 2
Note that both proofs are valid in the dynamic graph model mentioned in [16] since TVG is a stronger graph model. Therefore, agents need 1-hop visibility and global communication to solve the dispersion of $pn$ agents in the dynamic graph model mentioned in [16] as well.
Now, we provide the memory lower bound result to solve $k$-Dispersion on TVG. The proof of this lemma is motivated from [3]. In [3], they have provided the memory lower bound $\Omega(\log k)$ to solve dispersion on a static graph when $k\leq n$. We extend their idea for the memory lower bound when $k>n$.
Lemma 1
Assuming all agents are given the same amount of memory, agents require $\Omega(\log n)$ bits of memory each for any deterministic algorithm to solve $k$-Dispersion on TVG when $k\geq n+1$, and this proof is valid even if the adversary does not delete any edge at any round from footprint $G$.
Proof
Suppose all agents have $o(\log n)$ bits of memory. Each agent’s state space is $2^{o(\log n)}=n^{o(1)}$. Since there are at least $n+1$ agents, by pigeonhole principle, there exist at least $k-n+2$ agents with the same state space. Let agents $a_{1}$, $a_{2}$, …, $a_{k-n+2}$ have the same state space. Let us suppose that all agents are initially co-located at some node. Since all agents run the same deterministic algorithm, and agents $a_{1}$, $a_{2}$, …, $a_{k-n+2}$ are co-located initially, they will run the same moves. It means that agents $a_{1}$, $a_{2}$, …, $a_{k-n+2}$ remain stay together. Therefore, by the end of the algorithm, agents $a_{1}$, $a_{2}$, …, $a_{k-n+2}$ are at the same node. Solving dispersion requires a configuration where there is at least one agent at every node. We will never solve dispersion because there is no way for agents $a_{1}$, $a_{2}$, …, $a_{k-n+2}$ to settle down on at least two different nodes. Therefore, all $k$ agents are at most $n-1$ nodes by the end of the algorithm. Thus, to solve the dispersion problem, agents need $\Omega(\log n)$ memory each. This proof is for static graphs, and $TVG$ is a weaker model than static. Therefore, the same proof is also valid for TVGs. ∎
3 The Dispersion on General Graphs
In this section, we introduce an algorithm that solves the $(n+1)$-Dispersion on TVG when $n+1$ agents are co-located at some node in a general graph, and agents do not have global communication and 1-hop visibility, i.e., we assume that agents are equipped with zero-hop visibility and local communication. We use the idea of depth-first search (DFS) traversal by mobile agents that also solves the dispersion on static graphs [3]. For the sake of completeness, let us recall the idea of DFS traversal by a group of agents. At each node $v$, a group of agents chooses an unexplored port $p$ and traverses through it: (i) if it reaches a node $u$ which is an empty node (agents can understand this by the fact that there is no agent settled at $u$), then one agent settles at $u$, and stores port $p^{\prime}$ (through which it entered node $u$) as its parent port. Now, the unsettled agents try to move through port $(p^{\prime}+1)$ mod $\delta_{u}$ in the explore state, (ii) else, if the group of agents reach a node which already has a settled agent, the agent returns to node $v$ in the backtrack state. If the unsettled agents reach node $v$ in the backtrack state, then the agents compute $q=(p+1)$ mod $\delta_{v}$. If port $q$ is unexplored, it moves through port $q$ in the explore state. Else, if port $q$ equals the parent port of $v$, then the agents move through port $q$ in the backtrack state. This algorithm takes $4m$ rounds to run on any graph, where $m$ is the number of edges in that graph. The information required for the group of agents at each node to run this algorithm can be stored by the settled agent that is already present at that node.
Recall that, according to our model, an agent does not know if any associated edge w.r.t. its current position is missing or not. This makes the adversary powerful as it can precompute the agents’ decision at a node and accordingly can delete an associated edge at the start of that round. If the group of unsettled agents decides to move through port $p$ according to our algorithm in round $r$, the adversary can delete the edge corresponding to port $p$ in round $r$. Thus preventing the group of agents from moving in round $r$.
High-level idea: No group of agents moves indefinitely through which its movement was unsuccessful. Initially, there is only one group as the configuration is rooted. Whenever the group has an unsuccessful move, the agents divide into two groups, namely $G_{1}$ and $G_{2}$. These two groups run their dispersion algorithm via DFS separately. A missing edge can make the movement unsuccessful for both groups. On unsuccessful movement, $G_{1}$ never deviates from its DFS traversal path, whereas $G_{2}$ deviates after some finite waiting period. This keeps the algorithm running, and either $G_{2}$ gets dispersed, or $G_{1}$ can move one hop according to its DFS traversal. If $G_{2}$ gets dispersed, $G_{1}$ can understand it within some finite rounds and gets divided into two groups further. Else, if $G_{1}$ gets dispersed by moving one hop again and again, $G_{2}$ can understand it within some finite rounds and gets divided into two groups further. We ensure that none of $G_{1}$ (or $G_{2}$) divides further unless $G_{2}$ (or $G_{1}$) is already dispersed. This keeps the total number of groups bounded by 2. As the settled agents keep the information required for the DFS traversal of all the groups, it helps the agents store the information using $O(\log n)$ memory. As we have different algorithms for the groups and different criteria for the groups to be divided, while dividing a group into two, each new group is given a unique label so that $G_{1}\,(G_{2})$ can execute their respective algorithms.
3.1 Algorithm for general graph
In this section, we provide a detailed description of the algorithm for dispersion on a time-varying graph $\mathcal{G}$ with at most one dynamic edge. In our algorithm, unsettled agents may get divided into two groups, we call them $Group_{1}$, and $Group_{2}$ and denote them by $G_{1}$, and $G_{2}$ respectively. At any particular round, there can be no more than two groups. However, over time, if all the unsettled agents of one group get settled, the other group can be divided into two groups again. So, if we look throughout the run of the algorithm, there may exist several groups. To identify these groups, we provide unique labelling to each group whenever a group gets generated and store it in a variable. Now we provide the parameters maintained by each agent $a_{i}$.
  • $\bm{a_{i}.success}$: When an agent $a_{i}$ attempts to move from a node $u$ to $v$ via the edge $(u,\,v)$, in the round $t$, it is either successful or unsuccessful. If the agent reaches the node $v$ in the round $t+1$, it is a successful attempt, thus, $a_{i}.success=True$; otherwise, $a_{i}.success=False$. Each agent $a_{i}$ initially has $a_{i}.success=True$.
  • $\bm{a_{i}.skip}$: If agent $a_{i}$ skips a port, then it stores in this parameter. Initially, $a_{i}.skip=-1$
  • $\bm{a_{i}.grp\_label}$: It stores the ID of the group to which agent $a_{i}$ belongs. Initially $a_{i}.grp\_label=10$. If $a_{i}.grp\_label$ ends with 0, it implies $a_{i}$ is a part of $G_{1}$. Else, if $a_{i}.grp\_label$ ends with 1, it implies $a_{i}$ is a part of $G_{2}$. Later we elaborate on how this variable gets updated.
  • $\bm{a_{i}.count_{1}}$: If agent $a_{i}\in G_{1}$, then it stores the number of consecutive odd rounds with $a_{i}.success=False$.
  • $\bm{a_{i}.count_{2}}$: If agent $a_{i}\in G_{2}$ then it stores the number of consecutive odd rounds with $a_{i}.success=False$.
  • $\bm{a_{i}.count_{3}}$: If agent $a_{i}\in G_{2}$, then it stores the number of new DFS traversals it started from the time $a_{i}$ became a part of this group.
[…]

### v3

@@ 5.2 Impossibility with every node with at most $\lceil k/n\rceil$ agents
The problem of $k$-balanced dispersion is impossible to solve in temporally connected dynamic graphs. This holds even if nodes have infinite storage, and agents have infinite memory, full visibility, global communication, and complete knowledge of both $k$ and $n$.
6 Impossibility and lower bound for $\ell$-bounded 1-interval connected graphs
In this section, we provide impossibility results of the dispersion problem in $\ell$-bounded 1-interval connected graphs, where $\ell\geq 1$.
Figure 5: Movement of agents at round $r$ at nodes $u$, $u_{1}$, and $u_{2}$.
Theorem 6.1.
It is impossible to solve $k$-balanced dispersion in 1-bounded 1-interval connected graphs for any $n\geq 4$ if agents have only 1-hop visibility and f-2-f communication. The impossibility holds for all undispersed initial configurations of agents, even when nodes have infinite storage, agents have infinite memory and know the values of $n$ and $k$.
Proof.
Consider the ring of size $n$ as a footprint $G$. Let $k=pn$ agents be in an undispersed configuration in $G$, where $p\in\mathbb{N}$.vvvIf $k=pn$, then in a dispersed configuration, every node has $p$ agents. If the agents have 1-hop visibility, then as per our visibility model, the agent can sense the missing edge corresponding to some port $\lambda$. The movement of the agents depends on 1-hop visibility and f-2-f communication. We prove our result using a contrapositive argument. Suppose there is a deterministic algorithm $\mathcal{A}$ which solves dispersion.
Since the adversary knows the algorithm $\mathcal{A}$, it can precompute the agents’ movement in $\mathcal{G}_{r}$ for every round $r$. At round $r$, considering $\rho(e,r)=1$ for all $e\in E$, if it precomputes that it does not lead to a dispersed configuration, then it keeps $\mathcal{G}_{r}=G$ at the beginning of round $r$. Otherwise, it sets $\rho(e_{j},r)=0$ for a chosen edge $e_{j}\in E$, and $\rho(e,r)=1$ for every $e\in E\setminus\{e_{j}\}$. How the adversary selects such an edge $e_{j}$ at round $r$ is described below.
Since agents are not in a dispersed configuration at the beginning of round $r$, there is a node $u$ which has at most p1p-1 agents. Let $u_{1}$ and $u_{2}$ be two neighbours of node $u$ in $G$. Let $z$ ($\leq p-1$) agents be at node $u$. If $\mathcal{G}_{r}=G$, then the following is the precomputation of the agents as per algorithm $\mathcal{A}$: let $x_{1}$ agents move from node $u$ to $u_{1}$, $y_{1}$ agents move from node $u$ to $u_{2}$, $x_{2}$ agents move from node $u_{1}$ to $u$, and $y_{2}$ agents move from node $u_{2}$ to $u$ (see Fig. 5). And accordingly, at the end of round $r$, there are $z+(x_{2}-x_{1})+(y_{2}-y_{1})=p$ agents at node $u$ in $G$. The adversary knows the precomputed values of $x_{1},x_{2},y_{1},y_{2}$. It is important to note that $(x_{2}-x_{1})\leq 0$ and $(y_{2}-y_{1})\leq 0$ are not possible due to the following reason. If $(x_{2}-x_{1})\leq 0$ and $(y_{2}-y_{1})\leq 0$, then the value $z+(x_{2}-x_{1})+(y_{2}-y_{1})\leq z<p$ as $z<p$. Therefore, either $(x_{2}-x_{1})>0$ or $(y_{2}-y_{1})>0$. Without loss of generality, let $(y_{2}-y_{1})>0$. In this case, at the beginning of round $r$, the adversary constructs $\mathcal{G}_{r}$ as follows: $\rho(\overline{uu_{2}},r)=0$, and $\rho(e,r)=1$ for every $e\in E\setminus\{\overline{uu_{2}}\}$. Let’s call this new configuration $G^{\prime}$. Note that the 1-hop view of agents at node $u_{1}$ in $G^{\prime}$ is the same as the 1-hop view at node $u_{1}$ in $G$, and the 1-hop view of agents at node $u$ and $u_{2}$ changes at round $r$. Since f-2-f communication is present, agents at node $u_{1}$ can not get the information about the configuration change, and hence the movement of agents at node $u_{1}$ in $G^{\prime}$ remains the same as in $G$ at round $r$. Since the 1-hop view of agents at node $u$ and $u_{2}$ has changed, they can change their decision based on algorithm $\mathcal{A}$. In this case, if more than $x_{1}$ or fewer than $x_{1}$ agents move from $u$ to $u_{1}$ in $G^{\prime}$, then at node $u_{1}$, the number of agents is not $p$ due to the fact that the 1-hop view of node $u_{1}$ in $G^{\prime}$ is the same as in $G$, and $n\geq 4$. Therefore, the number of agents at node $u$ in $G^{\prime}$ is $z+(x_{2}-x_{1})$. If $z+(x_{2}-x_{1})=p$, then $z+(x_{2}-x_{1})=z+(x_{2}-x_{1})+(y_{2}-y_{1})\implies(y_{2}-y_{1})=0$. This gives a contradiction. Therefore, at the end of round $r$, the number of agents at node $u$ in $G^{\prime}$ is not $p$. Hence, the dispersion is not achieved in $G^{\prime}$ at the end of round $r$.
Since in every round $r\geq 0$, $\overline{E}_{r}\leq 1$, and in every round $\mathcal{G}_{r}$ is connected. Therefore, $\mathcal{G}$ is a 1-bounded 1-interval connected graph. Also, this proof is independent of nodes having infinite storage, agents having infinite memory and knowledge of $n$, $k$. This completes the proof. ∎
Theorem 6.2.
It is impossible to solve $k$-balanced dispersion in 1-bounded 1-interval connected graphs for any $n\geq 3$ if agents have only 0-hop visibility and global communication. The impossibility holds for all undispersed initial configurations of agents, even when agents have infinite memory and know the values of $n$ and $k$.
Proof.
Consider the ring of size $n$ as a footprint $G$. Let $k=pn$ agents be in an arbitrary configuration at nodes of $G$, where $p\in\mathbb{N}$. If the agents have 0-hop visibility, then the choice of movement of each agent depends only on the information present at the current node and the information attained via global communication. Therefore, the movement of the agents depends on 0-hop visibility and global communication. We prove our result using a contrapositive argument. Suppose there is a deterministic algorithm $\mathcal{A}$ which solves dispersion.
Since the adversary knows the algorithm $\mathcal{A}$, it can precompute the agents’ movement in $\mathcal{G}_{r}$ for every round $r$. At round $r$, considering $\rho(e,r)=1$ for all $e\in E$, if it precomputes that it does not lead to a dispersed configuration, then it keeps $\mathcal{G}_{r}=G$ at the beginning of round $r$. Otherwise, it sets $\rho(e_{j},r)=0$ for a chosen edge $e_{j}\in E$, and $\rho(e,r)=1$ for every $e\in E\setminus\{e_{j}\}$. How the adversary selects such an edge $e_{j}$ at round $r$ is described below.
Since agents are not in a dispersed configuration at the beginning of round $r$, there is a node $u$ which has at least $p+1$ agents. Let $u_{1}$ and $u_{2}$ be two neighbours of node $u$ in $G$. Let $z$ ($\geq p+1$) agents be at node $u$. If $\mathcal{G}_{r}=G$, then the following is the precomputation of the agents as per algorithm $\mathcal{A}$: let $x_{1}$ agents move from node $u$ to $u_{1}$, $y_{1}$ agents move from node $u$ to $u_{2}$, $x_{2}$ agents move from node $u_{1}$ to $u$, and $y_{2}$ agents move from node $u_{2}$ to $u$ (see Fig. 5). And accordingly, at the end of round $r$, there are $z+(x_{2}-x_{1})+(y_{2}-y_{1})=p$ agents at node $u$. The adversary knows the precomputed values of $x_{1},x_{2},y_{1},y_{2}$. It is important to note that $(x_{2}-x_{1})\geq 0$ and $(y_{2}-y_{1})\geq 0$ are not possible due to the following reason. If $(x_{2}-x_{1})\geq 0$ and $(y_{2}-y_{1})\geq 0$, the value of $z+(x_{2}-x_{1})+(y_{2}-y_{1})\geq z\geq p+1$. Therefore, either $(x_{2}-x_{1})<0$ or $(y_{2}-y_{1})<0$. Without loss of generality, let $(y_{2}-y_{1})<0$. In this case, at the beginning of round $r$, the adversary constructs $\mathcal{G}_{r}$ as follows: $\rho(\overline{uu_{2}},r)=0$, and $\rho(e,r)=1$ for every $e\in E\setminus\{\overline{uu_{2}}\}$. Let’s call this new configuration $G^{\prime}$. Since agents have 0-hop visibility, they can not understand the configuration change at round $r$. Hence, the movement of agents in $G^{\prime}$ is the same as in $G$. Since edge $\overline{uu_{2}}$ is not present, the movement of agents from node $u$ to $u_{2}$ and node $u_{2}$ to $u$ in $G^{\prime}$ will be unsuccessful. Therefore, the number of agents at node $u$ in $G^{\prime}$ is $z+(x_{2}-x_{1})$. If $z+(x_{2}-x_{1})=p$, then $z+(x_{2}-x_{1})=z+(x_{2}-x_{1})+(y_{2}-y_{1})\implies(y_{2}-y_{1})=0$. This gives a contradiction as $y_{2}<y_{1}$. Therefore, at the end of round $r$, the number of agents at node $u$ in $G^{\prime}$ is not $p$. Hence, the dispersion is not achieved in $G^{\prime}$ at the end of round $r$.
Since in every round $r\geq 0$, $\overline{E}_{r}\leq 1$, and in every round $\mathcal{G}_{r}$ is connected. Therefore, $\mathcal{G}$ is a 1-bounded 1-interval connected graph. Also, this proof is independent of nodes having infinite storage, agents having infinite memory and knowledge of $n$, $k$. This completes the proof. ∎
Remark 6.1.
Theorems 6.1 and 6.2 hold when $k=pn$, for any $p\in\mathbb{N}$. That is, a team of $k=pn$ agents requires both 1-hop visibility and 1-hop communication to solve $k$-balanced dispersion in 1-bounded 1-interval connected graphs.
Remark 6.2.
Theorems 6.1 and 6.2 are valid in $\ell$-bounded 1-interval connected graphs for any $\ell\geq 1$. In $\ell$-bounded 1-interval connected graphs, solving $k$-balanced dispersion with $k=pn$ agents (where $p\in\mathbb{N}$) also requires both 1-hop visibility and 1-hop communication.
Now, we show two impossibilities to solve $k$-balanced dispersion in $\ell$-bounded 1-interval connected graphs.
Theorem 6.3.
($\ell\geq 25$) It is impossible to solve $k$-balanced dispersion in $\ell$-bounded 1-interval connected graphs when agents have $1$-hop visibility and f-2-f communication. The impossibility holds for all initial configurations where at least one node has more than $\lceil k/n\rceil$ agents, even if nodes have infinite storage, agents have infinite memory and knowledge of $n$, $\ell$, and $k$.
Proof.
In this proof, we consider $\ell\geq 25$ and the clique of size $n$, where $n\in\big{[}5,\lfloor\sqrt{l}\rfloor\big{]}$, as a footprint $G$. Consider $k=pn+q$, where $p,q\in\mathbb{N}\cup\{0\}$ and $5\leq q\leq n-1$. Consider an initial configuration which has at least one node containing at least $\lceil k/n\rceil+1=p+2$ agents. Let $S_{\mathcal{G}_{r}}(u)$ denote the number of agents at node $u$ at the beginning of round $r$, and $E_{\mathcal{G}_{r}}(u)$ denote the number of agents at node $u$ at the end of round $r$. At the beginning of round $r\geq 0$, the adversary finds a path $w_{1}\sim w_{2}\sim\ldots\sim w_{n}$ of length $n$ in $G$ such that $S_{\mathcal{G}_{r}}(w_{i})\geq S_{\mathcal{G}_{r}}(w_{i+1})$ holds for every $i\in[1,n-1]$. Since $G$ is a clique, the adversary can always find at least one such path in $G$ at the beginning of round $r$. If there are several such paths, the adversary can pick any. Let’s call the path chosen by the adversary as $P_{r}$. Let $E^{\prime}$ be the edges of $P_{r}$. At the beginning of round $r$, the adversary sets $\mathcal{G}_{r}$ to either $P_{r}$ or $P_{r}^{\prime}$, where $P_{r}^{\prime}$ is defined later.
Suppose there is an algorithm $\mathcal{A}$ that solves dispersion. Since the adversary is aware of $\mathcal{G}_{r}$ and the agent’s algorithm, it can precompute the movement of agents at round $r$. Based on precomputation, if the adversary observes that at the end or round $r$, there is at least one node in $P_{r}$ which has at least $p+2$ agents, then it constructs $\mathcal{G}_{r}=P_{r}$ at the beginning of round $r$. Otherwise, based on precomputation, if the adversary observes that at the end or round $r$, there is no node in $P_{r}$ which has at least $p+2$ agents, it sets $\mathcal{G}_{r}$ to $P_{r}^{\prime}$ such that there is at least one node in $P_{r}^{\prime}$ with at least $p+2$ agents at the end of round $r$. Let $E^{\prime\prime}=\big{(}E^{\prime}\setminus\{\overline{w_{4}w_{5}}\}\big{)}\cup\{\overline{w_{1}w_{5}}\}$. The graph $P_{r}^{\prime\prime}$ is $(V,E^{\prime\prime})$.
Figure 6: Agents’ movement when $\mathcal{G}_{r}$ is $P_{r}$.
Observe that in $P_{r}^{\prime}$, the 1-hop view of agents located at nodes $w_{1}$, $w_{4}$, and $w_{5}$ is changed, while the 1-hop view of agents at all other nodes remains identical to their view in $P_{r}$, due to the presence of f-2-f communication. Now we prove that there is a node in $P_{r}^{\prime}$ which has at least $p+2$ agents. Since there is a node with at least $p+2$ agents in $\mathcal{G}_{r}$ at the beginning of round $r$, and $S_{\mathcal{G}_{r}}(w_{i})\geq S_{\mathcal{G}_{r}}(w_{i+1})$ for every $i$, therefore, $S_{P_{r}}(w_{1})\geq p+2$. As per precomputaion in $P_{r}$, let $x_{i}$ agents move from node $w_{i}$ to $w_{i+1}$, and $y_{i+1}$ agents move from node $w_{i+1}$ to $w_{i}$ (see Fig. 6 for agents’ movement in $P_{r}$). Since at the end of round $r$, there is no node with at least $p+2$ agents in $P_{r}$, the following cases are possible.
  • 1.
    Case 1: $E_{P_{r}}(w_{3})=p-c_{1}$ and $E_{P_{r}}(w_{4})=p-c_{2}$, where $c_{1}\geq 0$ and $c_{2}\geq 0$,
  • 2.
    Case 2: $E_{P_{r}}(w_{3})=p+1$ and $E_{P_{r}}(w_{4})=p-c_{2}$, where $c_{2}\geq 0$,
  • 3.
    Case 3: $E_{P_{r}}(w_{3})=p-c_{1}$ and $E_{P_{r}}(w_{4})=p+1$, where $c_{1}\geq 0$,
  • 4.
    Case 4: $E_{P_{r}}(w_{3})=p+1$ and $E_{P_{r}}(w_{4})=p+1$.
Since $q\geq 5$, $S_{P_{r}}(w_{1})\geq p+2$ and $S_{P_{r}}(w_{i})\geq S_{P_{r}}(w_{i+1})$ in $P_{r}$, the value of $\sum_{i=1}^{4}S_{P_{r}}(w_{i})$ is at least $4p+5$. Hence,
$x_{4}-y_{5}=\sum_{i=1}^{4}S_{P_{r}}(w_{i})-\sum_{i=1}^{4}E_{P_{r}}(w_{i})\geq 4p+5-\sum_{i=1}^{4}E_{P_{r}}(w_{i}).$
Since at the end of round $r$, there is no node in $P_{r}$ which has at least $p+2$ agents, $E_{P_{r}}(w_{1})\leq p+1$ and $E_{P_{r}}(w_{2})\leq p+1$. Hence,
$x_{4}-y_{5}\geq 2p+3-\sum_{i=3}^{4}E_{P_{r}}(w_{i}).$
Since in all aforementioned cases, we have the value of $E_{P_{r}}(w_{3})$ and $E_{P_{r}}(w_{4})$, the value of $x_{4}-y_{5}$ is:
  • 1.
    Case 1 $\implies$ $x_{4}-y_{5}\geq 3+c_{1}+c_{2}$
  • 2.
    Case 2 $\implies$ $x_{4}-y_{5}\geq 2+c_{2}$,
  • 3.
    Case 3 $\implies$ $x_{4}-y_{5}\geq 2+c_{1}$,
  • 4.
    Case 4 $\implies$ $x_{4}-y_{5}\geq 1$.
Figure 7: Agents’ movement when $\mathcal{G}_{r}$ is $G^{\prime}$.
Note that if $\mathcal{G}_{r}=P_{r}$, then $E_{P_{r}}(w_{3})=S_{P_{r}}(w_{3})+(y_{4}+x_{2})-(y_{3}+x_{3})$, and $E_{P_{r}}(w_{4})=S_{P_{r}}(w_{4})+(y_{5}+x_{3})-(y_{4}+x_{4})$. Since the 1-hop view of node $w_{2}$ and $w_{3}$ in $P_{r}^{\prime}$ is the same as in $P_{r}$, the decision of agents’ movement at node $w_{2}$ and $w_{3}$ in $P_{r}^{\prime}$ remains the same as in $P_{r}$. Note that in $P_{r}^{\prime}$, $S_{P_{r}}(w_{i})=S_{P_{r}^{\prime}}(w_{i})$ for every $i$. Since there is no edge between node $w_{4}$ and $w_{5}$, no agent can move from node $w_{4}$ to $w_{5}$ in $P_{r}^{\prime}$. Therefore, in $P_{r}^{\prime}$, $E_{P_{r}^{\prime}}(w_{4})=S_{P_{r}^{\prime}}(w_{4})+x_{3}-y_{4}^{\prime}$ and $E_{P_{r}^{\prime}}(w_{3})=S_{P_{r}^{\prime}}(w_{3})+(y_{4}^{\prime}+x_{2})-(y_{3}+x_{3})$, where $y_{4}^{\prime}$ is the number agents that move from node $w_{4}$ to node $w_{3}$ in $P_{r}^{\prime}$ as the 1-hop view of $w_{2}$, $w_{3}$ are the same but the 1-hop view of node $w_{4}$ is changed (See Fig. 7 for agents’ movement in $P_{r}^{\prime}$). Observe that if $x_{4}-x_{5}\leq 0$, then the value of $\sum_{i=1}^{4}E_{P_{r}}(w_{i})$ is at least $\sum_{i=1}^{4}S_{P_{r}}(w_{i})$. Therefore, $\sum_{i=1}^{4}E_{P_{r}}(w_{i})\geq 4p+5$ as $\sum_{i=1}^{4}S_{P_{r}}(w_{i})\geq 4p+5$. This contradicts the fact that at the end of round $r$, at most $p+1$ agents are present at every node in $P_{r}$. So, the value of $x_{4}-x_{5}$ must be at least 1. In $P_{r}^{\prime}$, one might argue that if $x_{1}^{\prime\prime}-y_{5}^{\prime}=x_{4}-y_{5}$ holds, where $x_{1}^{\prime\prime}$ agents move from node $w_{1}$ to $w_{5}$ in $P_{r}^{\prime}$, and $y_{5}^{\prime}$ agents move from node $w_{5}$ to $w_{1}$ in $P_{r}^{\prime}$, then the extra load of agents can be distributed across nodes $w_{5},w_{6},\ldots,w_{n}$, and it might be possible to ensure that at the end of round $r$, at most $p+1$ agents are present at every node in $P_{r}^{\prime}$. Note that agents at nodes $w_{1}$, $w_{4}$, and $w_{5}$ are aware of this change. However, agents at nodes $w_{2}$ and $w_{3}$ are not aware of it; as previously mentioned, the decisions of agents at $w_{2}$ and $w_{3}$ remain the same as in $P_{r}$. Below, we show that due to this reason, there is a node with at least $p+2$ agents at the end of round $r$ in $P_{r}^{\prime}$.
  • 1.
    Case 1: Here, $x_{4}-y_{5}\geq 3+c_{1}+c_{2}$. We divide this case into the following sub-cases.
    • (a)
      If $y_{4}^{\prime}\leq y_{4}+1+c_{1}$,
      $E_{P_{r}^{\prime}}(w_{4})$ $=S_{P_{r}^{\prime}}(w_{4})+x_{3}-y_{4}^{\prime}$
[…]
      $E_{P_{r}^{\prime}}(w_{4})$ $\geq S_{P_{r}}(w_{4})+x_{3}-y_{4}-1-c_{1},\;\text{ as }\;y_{4}^{\prime}\leq y_{4}+c_{1}+1$
      $=S_{P_{r}}(w_{4})+x_{3}-y_{4}+(y_{5}-x_{4})-(y_{5}-x_{4})-1-c_{1}$
      $=S_{P_{r}}(w_{4})+(y_{5}+x_{3})-(y_{4}+x_{4})+(x_{4}-y_{5})-1-c_{1}$
      $=p+(x_{4}-y_{5})-1-c_{1}-c_{2},\;\;$
       $\;\;\text{ as }\;S_{P_{r}}(w_{4})+(y_{5}+x_{3})-(y_{4}+x_{4})=E_{P_{r}}(w_{4})=p-c_{2}$
      $\therefore E_{P_{r}^{\prime}}(w_{4})$ $\geq p+2,\;\text{ as }\;x_{4}-y_{5}\geq 3+c_{1}+c_{2}$
    • (b)
      If $y_{4}^{\prime}\geq y_{4}+2+c_{1}$,
      $E_{P_{r}^{\prime}}(w_{3})$ $=S_{P_{r}^{\prime}}(w_{3})+(y_{4}^{\prime}+x_{2})-(y_{3}+x_{3})$
[…]
      $E_{P_{r}^{\prime}}(w_{3})$ $\geq S_{P_{r}}(w_{3})+(y_{4}+x_{2})-(y_{3}+x_{3})+c_{1}+2,\;\text{ as }\;y_{4}^{\prime}\geq y_{4}+c_{1}+2$
      $=p+2,\;\text{ as }\;S_{P_{r}}(w_{3})+(y_{4}+x_{2})-(y_{3}+x_{3})=E_{P_{r}}(w_{3})=p-c_{1}$
[…]